# DRtool: An Interactive Tool for Analyzing High-Dimensional Clusterings

### v2

@@ 4.1 Image Data Example
To demonstrate use of the tool, we explore the MNIST data set in detail. Again, the $784\times 784$-pixel images were flattened and 1,500 samples were randomly sampled. A PCA pre-processing step was applied prior to applying UMAP (McInnes et al., 2018) to construct a two-dimensional embedding. To replicate a real use case, we study a k-means clustering instead of the true class labels (Figure 3). The reader may follow along using the run_example(example="MNIST", cluster="kmeans") function in our DRtools package.
@@ 4.1.1 Classes 1 and 2
There seems to be minimal separation between classes 1 and 2, suggesting they may correspond to the same digit. We select a path from point 25,483 in class 1 to point 44,483 in class 2. To get a closer look, we first look at the Path Projection Plot. The chosen number of dimensions is 100, which retains 97% of the variance, and the path stabilizes at a CCA degree of 4.
The resulting plot depicts overlap between the two classes. Adjusting the bandwidth of the density estimate to 1.5 shows unimodal density, suggesting the two classes may come from the same population. Showing the MST edges also does not provide any evidence of separation. The MST test results, however, may suggest otherwise. Seven crossings are counted when the approximate expectation under the null is 11.03 with a standard error of 3.523. While the bootstrapped $p$-value of $0.06$ is insignificant at the 5% level, the closeness indicates a more careful examination is necessary.
Inspection of the handwritten digits themselves reveals an interesting trend. While the majority of samples from both classes depict the digit one, the angle of the stroke differs drastically between the two classes (Figure 4). Following the path from class 1 to class 2, the strokes become more slanted. Both the MST path and MST test were able to detect this phenomenon, even though the two classes technically corresponded to the same digit.
@@ 4.1.2 Class 4
Class 4 is split between two different clusters in the UMAP embedding. We use the drawing tool to select the two clusters as our groups. The path projection settings are calibrated to 100 dimensions and a CCA degree of three. We also select the Group Coloring setting, so the points are colored according to group, rather than class. Analysis of the plot and estimated density does not provide evidence of separation. The MST edges, however, are more revealing after close inspection. There are few inter-group edges, even in overlapping regions (Figure 5). On the contrary, the MST test counts a larger number of edge crossings. 14 are counted when only 12.02 are expected with a standard error of 3.378.
@@ 4.2 Mass Cytometry Data Set
We now explore a mass cytometry data set (Wong et al., 2016) covering 35 samples originating from eight distinct human tissues enriched for T and natural killer cells. The data is processed and labeled inline with the procedure used in Becht et al. (2019). 3,000 cells were randomly sampled. To replicate a real use case, we explore a $k$-means clustering instead of the true class labels (Figure 7). The reader may follow along using the run_example(example="Wong", cluster="kmeans") function.
@@ 4.2.1 Class 4
To better understand why these two clusters are separated despite minimal evidence of separation, we reference the heatmap and meta data. According to the heatmap, the two groups differ most in CD8 T cell counts. This is confirmed by the cell labels provided by Becht et al. (2019), which were passed to the tool as meta data. So while separation wasn’t observed by the MST, the discrepancy in CD8 T cell counts accounts for the splitting of the class 4 points.
@@ 4.3 Class 8
Running the MST test returns interesting results. 20 crossings are counted when only 16.11 are expected with a standard error 4.44. The number of crossings are well above the number expected under the null hypothesis. This is due to the disparity in group densities. In order to preserve the size of the test, the null distribution must be constructed using a unimodal distribution whose density is similar to the that of the lesser-dense cluster (Section 2.4.2). This is why the expected number of crossings is so low. This phenomenon, however, has statistical meaning. The low relative density of group 1 provides minimal evidence from which we may draw conclusions. This lack of confidence is reflected in the low null expectation, and thus, the inflated p-value.
Overall, there is not enough evidence to declare the two groups are correctly separated. This it not quite correct according to the true labels. The two groups were not sampled from the same type of tissue, but the situation is slightly more complicated. See Figure B2 for details.
4.4 Class 3
The class 3 points are separated into three clusters – two larger, elongated clusters to the left, and one smaller cluster to the right (Figure 7). Initial analysis of the path projection plot and MST test did not reveal any evidence of separation. However, the heatmap and meta data explain the separation captured by UMAP. The two larger clusters to the left are completely disjoint in their CD161 gene counts, but very similar in all other features. Meanwhile, the third smaller cluster is distinct from the other two larger clusters in its TCRgD counts. The meta data also reveals the left two clusters consist of CD8 T cells, while the right cluster consists of Tgd cells.
@@ 5 Discussion
We have introduced our R package, DRtool, and exemplified its use cases. The MST serves as an effective medium for understanding high-dimensional relationships and structures. The various analytical tools provided by the package allow the user to extract a maximal amount of information from the MST by providing multiple prospectives. Such a multi-faceted view is necessary to understand contemporary dimension reduction methods that are trying to fit hundreds, or even thousands, of dimensions-worth of information into only two dimensions. Advances in multiple fields have lead to a surge in complex data, necessitating tools such as ours that help analysts assess and confirm their dimension reduction results.
@@ 6 Code Availability
All data and code a freely available at https://wwww.github.com/JustinMLin/DRtool.
@@ 7 Declaration of Interest
Appendices
Appendix A Details for the Null Hypothesis Problem
Let $n_{1},n_{2}>0$ be the sample sizes of each group and $c\in[-1,1]$ the location of the mode. Without loss of generality, assume $c\geq 0$. Let $\mathcal{F}$ be the family of distributions on [1,1][-1,1] such that
  • $f$ is increasing on [1,c][-1,c],
  • $f$ is decreasing on $[c,1]$,
  • $\int_{-1}^{0}f=\frac{n_{1}}{n_{1}+n_{2}}$, and
  • $\int_{0}^{1}f=\frac{n_{2}}{n_{1}+n_{2}}$.
A.1 Case I. $c>\epsilon$
Lemma A.1.
If $c\geq 0$ and $\epsilon>0$, then
$\int_{-\epsilon}^{0}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}}$
for all $f\in\mathcal{F}$.
Proof.
Let $f\in\mathcal{F}$. By means of contradiction, suppose
$\int_{-\epsilon}^{0}f<\epsilon\frac{n_{1}}{n_{1}+n_{2}}.$
Then
$\int_{-1}^{-\epsilon}f=\int_{-1}^{0}f-\int_{-\epsilon}^{0}f=\frac{n_{1}}{n_{1}+n_{2}}-\int_{-\epsilon}^{0}f>(1-\epsilon)\frac{n_{1}}{n_{1}+n_{2}}.$
Hence, there exists $x_{0}\in(-1,-\epsilon)$ such that
$f(x_{0})>\frac{n_{1}}{n_{1}+n_{2}}.$
By the unimodality constraint,
$f(x)\geq f(x_{0})>\frac{n_{1}}{n_{1}+n_{2}}\textrm{ for all }x\in[-\epsilon,0].$
This implies
$\int_{-\epsilon}^{0}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}},$
a contradiction. Therefore,
$\int_{-\epsilon}^{0}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}}.$
Let $f\in\mathcal{F}$. By Lemma A.1,
$f(x)\geq\sup\{f(x):-\epsilon\leq x\leq 0\}\geq\frac{n_{1}}{n_{1}+n_{2}}\textrm{ for all }x\in[0,\epsilon).$
Hence,
$\int_{0}^{\epsilon}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}}.$
We’ve shown
$\int_{-\epsilon}^{\epsilon}f=\int_{-\epsilon}^{0}f+\int_{0}^{\epsilon}f\geq 2\epsilon\frac{n_{1}}{n_{1}+n_{2}}\textrm{ for all }f\in\mathcal{F}.$
Therefore, the function $f^{\prime}:[-1,1]\to\mathbb{R}$ defined by
$f^{\prime}(x)=\begin{cases}\frac{n_{1}}{n_{1}+n_{2}}&-1\leq x\leq\epsilon\\ \frac{n_{2}-\epsilon n_{1}}{(1-\epsilon)(n_{1}+n_{2})}&\epsilon<x\leq 1\end{cases}$
belongs to $\mathcal{F}$ and minimizes $\int_{-\epsilon}^{\epsilon}f$.
A.2 Case II. $0<c<\epsilon$ and $\frac{n_{2}}{n_{1}}\geq c$
Let $f\in\mathcal{F}$. By Lemma A.1, there exists $x_{0}\in(-\epsilon,0]$ such that $f(x_{0})\geq\frac{n_{1}}{n_{1}+n_{2}}$. Thus,
$f(x)\geq f(x_{0})\geq\frac{n_{1}}{n_{1}+n_{2}}\textrm{ for all }x\in[0,c)$
$\Rightarrow\int_{0}^{c}f\geq c\frac{n_{1}}{n_{1}+n_{2}}.$
By the unimodality constraint,
$\frac{\int_{c}^{\epsilon}f}{\int_{\epsilon}^{1}f}\geq\frac{\epsilon-c}{1-\epsilon}$
$\begin{split}\Rightarrow\int_{c}^{\epsilon}f&\geq\frac{\epsilon-c}{1-\epsilon}\int_{\epsilon}^{1}f\\ &=\frac{\epsilon-c}{1-\epsilon}\left(\int_{0}^{1}f-\int_{0}^{c}f-\int_{c}^{\epsilon}f\right)\\ &=\frac{\epsilon-c}{1-\epsilon}\frac{n_{2}}{n_{1}+n_{2}}-\frac{\epsilon-c}{1-\epsilon}\int_{0}^{c}f-\frac{\epsilon-c}{1-\epsilon}\int_{c}^{\epsilon}f\end{split}$
$\Rightarrow\frac{1-c}{1-\epsilon}\int_{c}^{\epsilon}f\geq\frac{\epsilon-c}{1-\epsilon}\frac{n_{2}}{n_{1}+n_{2}}-\frac{\epsilon-c}{1-\epsilon}\int_{0}^{c}f$
$\Rightarrow\int_{c}^{\epsilon}f\geq\frac{\epsilon-c}{1-c}\frac{n_{2}}{n_{1}+n_{2}}-\frac{\epsilon-c}{1-c}\int_{0}^{c}f$
It follows that
$\begin{split}\int_{0}^{\epsilon}f&=\int_{0}^{c}f+\int_{c}^{\epsilon}f\\ &\geq\int_{0}^{c}f+\frac{\epsilon-c}{1-c}\frac{n_{2}}{n_{1}+n_{2}}-\frac{\epsilon-c}{1-c}\int_{0}^{c}f\\ &=\left(1-\frac{\epsilon-c}{1-c}\right)\int_{0}^{c}f+\frac{\epsilon-c}{1-c}\frac{n_{2}}{n_{1}+n_{2}}\\ &\geq\left(1-\frac{\epsilon-c}{1-c}\right)c\frac{n_{1}}{n_{1}+n_{2}}+\frac{\epsilon-c}{1-c}\frac{n_{2}}{n_{1}+n_{2}}\\ &=c\frac{n_{1}}{n_{1}+n_{2}}+\frac{(\epsilon-c)(n_{2}-cn_{1})}{(1-c)(n_{1}+n_{2})}.\end{split}$
Hence,
$\int_{-\epsilon}^{\epsilon}f=\int_{-\epsilon}^{0}f+\int_{0}^{\epsilon}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}}+c\frac{n_{1}}{n_{1}+n_{2}}+\frac{(\epsilon-c)(n_{2}-cn_{1})}{(1-c)(n_{1}+n_{2})}.$
Therefore, the function $f^{\prime}:[-1,1]\to\mathbb{R}$ defined by
$f^{\prime}(x)=\begin{cases}\frac{n_{1}}{n_{1}+n_{2}}&-1\leq x\leq c\\ \frac{n_{2}-cn_{1}}{(1-c)(n_{1}+n_{2})}&c<x\leq 1\end{cases}$
belongs to $\mathcal{F}$ and minimizes $\int_{-\epsilon}^{\epsilon}f$.
A.3 Case III. $0<c<\epsilon$ and $\frac{n_{2}}{n_{1}}<c$
Let $f\in\mathcal{F}$. First we show
$\int_{0}^{\epsilon}f\geq\epsilon\frac{n_{2}}{n_{1}+n_{2}}.$
By means of contradiction, assume
$\int_{0}^{\epsilon}f<\epsilon\frac{n_{2}}{n_{1}+n_{2}}.$
There exists $x_{0}\in[0,\epsilon)$ such that
$f(x_{0})<\frac{n_{2}}{n_{1}+n_{2}}.$
If $x_{0}\leq c$, then
$f(x)\leq f(x_{0})<\frac{n_{2}}{n_{1}+n_{2}}<\frac{n_{1}}{n_{1}+n_{2}}\textrm{ for all }x\in(-\epsilon,0]$
$\Rightarrow\int_{-\epsilon}^{0}f<\epsilon\frac{n_{1}}{n_{1}+n_{2}},$
contradicting Lemma A.1.
If $x_{0}>c$, then
$f(x)\leq f(x_{0})<\frac{n_{2}}{n_{1}+n_{2}}\textrm{ for all }x\in[\epsilon,1]$
$\Rightarrow\int_{\epsilon}^{1}f<(1-\epsilon)\frac{n_{2}}{n_{1}+n_{2}}.$
Thus,
$\int_{0}^{1}f=\int_{0}^{\epsilon}f+\int_{\epsilon}^{1}f<\epsilon\frac{n_{2}}{n_{1}+n_{2}}+(1-\epsilon)\frac{n_{2}}{n_{1}+n_{2}}=\frac{n_{2}}{n_{1}+n_{2}},$
contradicting $f\in\mathcal{F}$. Therefore,
$\int_{0}^{\epsilon}f\geq\epsilon\frac{n_{2}}{n_{1}+n_{2}}.$
We’ve shown
$\int_{-\epsilon}^{\epsilon}f=\int_{-\epsilon}^{0}f+\int_{0}^{\epsilon}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}}+\epsilon\frac{n_{2}}{n_{1}+n_{2}}.$
Therefore, the function $f^{\prime}:[-1,1]\to\mathbb{R}$ defined by
$f^{\prime}(x)=\begin{cases}\frac{n_{1}}{n_{1}+n_{2}}&-1\leq x\leq 0\\ \frac{n_{2}}{n_{1}+n_{2}}&0<x\leq 1\end{cases}$
belongs to $\mathcal{F}$ and minimizes $\int_{-\epsilon}^{\epsilon}f$.
A.4 Case IV. $c=0$
Let $f\in\mathcal{F}$. By Lemma A.1,
$\int_{-\epsilon}^{0}f\geq\epsilon\frac{n_{1}}{n_{1}+n_{2}}.$
A symmetrical argument shows
$\int_{0}^{\epsilon}f\geq\epsilon\frac{n_{2}}{n_{1}+n_{2}}.$
Therefore, the function $f^{\prime}:[-1,1]\to\mathbb{R}$ defined by
$f^{\prime}(x)=\begin{cases}\frac{n_{1}}{n_{1}+n_{2}}&-1\leq x\leq 0\\ \frac{n_{2}}{n_{1}+n_{2}}&0<x\leq 1\end{cases}$
belongs to $\mathcal{F}$ and minimizes $\int_{-\epsilon}^{\epsilon}f$.
Appendix B Supplementary Figures and Tables
Figure B1: MNIST Embedding with True Labels
Figure B2: Wong Embedding with True Labels
Table B1: MST Test Results with Increased Sample Size (Digits 3,5,8)

### v3

@@ 5.1 Image Data Example
To demonstrate use of the tool, we explore the MNIST data set in detail. The $(28\times 28)$-pixel images were flattened and 1,500 samples were randomly sampled. A PCA pre-processing step was applied prior to applying UMAP (UMAP) to construct a two-dimensional embedding. To replicate a real use case, we study a k-means clustering, calculated on the high-dimensional data, instead of the true class labels (Figure 5). The reader may follow along using the run_example(example="MNIST", cluster="kmeans") function in our DRtools package.
@@ 5.1.1 Classes 1 and 2
There seems to be minimal separation between classes 1 and 2, suggesting they may correspond to the same digit. We select a path from point 25,483 in class 1 to point 44,483 in class 2 (Figure 6). To get a closer look, we first look at the Path Projection Plot. The chosen number of dimensions is 100, which retains 97% of the variance, and the path stabilizes at a CCA degree of four.
Figure 6: Path between class 1 and class 2 in UMAP embedding of MNIST data set.
The resulting plot depicts overlap between the two classes (Figure 7). Adjusting the bandwidth of the density estimate to 1.5 shows unimodal density, suggesting the two classes may come from the same population. Showing the MST edges also does not provide any evidence of separation. The MST test results, however, may suggest otherwise. Seven crossings are counted when the approximate expectation under the null is 11.03 with a standard error of 3.523. While the bootstrapped $p$-value of $0.06$ is insignificant at the 5% level, the closeness indicates a more careful examination is necessary.
@@
Inspection of the handwritten digits themselves reveals an interesting trend. While the majority of samples from both classes depict the digit one, the angle of the stroke differs drastically between the two classes. Following the path from class 1 to class 2, the strokes become more slanted. Both the MST path and MST test were able to detect this phenomenon, even though the two classes technically corresponded to the same digit.
@@ 5.1.2 Class 4
Class 4 is split between two different clusters in the UMAP embedding. We use the drawing tool to select the two clusters as our groups. The path projection settings are calibrated to 100 dimensions and a CCA degree of three. We also select the Group Coloring setting, so the points are colored according to group, rather than class. Analysis of the plot and estimated density does not provide evidence of separation. The MST edges, however, are more revealing after close inspection. If you look closely, there are few inter-group edges, even in overlapping regions (Figure 8). On the contrary, the MST test counts a larger number of edge crossings. 14 are counted when only 12.02 are expected with a standard error of 3.378.
@@ 5.2 Mass Cytometry Data Set
We now explore a mass cytometry data set (Wong_data_set) covering 35 samples originating from eight distinct human tissues enriched for T and natural killer cells. The data is processed and labeled inline with the procedure used in UMAP_example. 3,000 cells were randomly sampled. To replicate a real use case, we explore a $k$-means clustering, calculated on the high-dimensional data, instead of the true class labels (Figure 10). The reader may follow along using the run_example(example="Wong", cluster="kmeans") function.
@@ 5.2.1 Class 4
To better understand why these two clusters are separated despite minimal evidence of separation, we reference the heatmap and meta data. According to the heatmap (Figure 12), the two groups differ most in CD8 protein counts. This is confirmed by the cell labels provided by UMAP_example, which were passed to the tool as meta data (Figure 13). So while separation wasn’t observed by the MST, the discrepancy in CD8 protein counts accounts for the splitting of the class 4 points.
Figure 12: Heatmap of cell counts within the two Class 4 sub-clusters.
Figure 13: Distribution of cell labels among cells in the two class 4 sub-clusters.
@@ 5.2.2 Class 8
Running the MST test returns interesting results. 20 crossings are counted when only 16.11 are expected with a standard error 4.44. The number of crossings is greater than the number expected under the null hypothesis. This is due to the disparity in group densities. In order to preserve the size of the test, the null distribution must be constructed using a unimodal distribution whose density is similar to the that of the lesser-dense cluster (Section 3.4.2). This is why the expected number of crossings is so low. This phenomenon, however, has statistical meaning. The low relative density of group 1 provides minimal evidence from which we may draw conclusions. This lack of confidence is reflected in the low null expectation, and thus, the inflated p-value.
Overall, there is not enough evidence to declare the two groups are correctly separated. This is not quite correct according to the true labels. The two groups were not sampled from the same type of tissue, but the situation is slightly more complicated. See Figure B2 for details.
5.2.3 Class 3
The class 3 points are separated into three clusters – two larger, elongated clusters to the left, and one smaller cluster to the right. Initial analysis of the path projection plot and MST test did not reveal any evidence of separation. However, the heatmap and meta data explain the separation captured by UMAP. The two larger clusters to the left are completely disjoint in their CD161 gene counts, but very similar in all other features. Meanwhile, the third smaller cluster is distinct from the other two larger clusters in its TCRgD counts. The meta data also reveal the left two clusters consist of CD8 T cells, while the right cluster consists of Tgd cells.
@@ 6 Discussion
We have introduced our R package, DRtool, and exemplified its use cases. The MST serves as an effective medium for understanding high-dimensional relationships and structures. The various analytical tools provided by the package allow the user to extract the maximal amount of information from the MST by providing multiple perspectives. Such a multi-faceted view is necessary to understand contemporary dimension reduction methods that are trying to fit hundreds, or even thousands, of dimensions-worth of information into only two dimensions. Advances in multiple fields have lead to a surge in complex data, necessitating tools such as ours that help analysts assess and confirm their high-dimensional clustering results.
@@ 7 Code Availability
All data and code are freely available at https://wwww.github.com/JustinMLin/DRtool.